what is the linear equation of a line with a slope of -2/7 that passes through the point (8,-7)
Answer by vksarvepalli(154)

equation of a line with slope m and passing through the point (x1,y1) is given by

(y-y1) = m(x-x1)
so equation of a line with a slope of -2/7 passing through the point (8,-7) is given by
(y-(-7))=(-2/7)(x-8)
so (y+7)=(-2/7)(x-8)
7y+49 = -2x+16
2x+7y+33=0